# FedStrategist: A Meta-Learning Framework for Adaptive and Robust Aggregation in Federated Learning

### v1

@@ 7.1 The Brittleness of Static Defenses and the Need for Adaptation
7.2 Interpreting the “Paradox of the Losing Winner”
@@
Our adaptive agent, in contrast, behaved as an effective risk manager. Its diagnostic metrics correctly identified a persistent attack, and it consequently learned a conservative policy that prioritized model integrity over raw, compromised performance. The slightly lower accuracy of the adaptive agent is not a failure, but rather the measurable “cost of security” paid to produce a reliable final model. This highlights that the objective in secure FL is not merely accuracy, but trustworthy accuracy.
@@ 7.3 The Controllable Risk Posture of FedStrategist
This demonstrates that FedStrategist is not a black box but a tunable controller. It reframes the defense problem from a search for a single “best” algorithm to providing a system where a practitioner can define their desired security posture via a single, interpretable parameter. This tunability is a critical feature for deploying FL systems in real-world applications with varying risk requirements.
@@ 8 Conclusion
Most significantly, our analysis of the stealth attack scenario revealed a crucial insight: optimizing for raw test accuracy alone is insufficient and can be misleading. We showed that FedStrategist successfully navigates the ”paradox of the losing winner,” learning a conservative policy to guarantee model integrity even when a compromised, non-robust baseline appeared to perform better. Furthermore, we demonstrated that our framework is not a black box, but a controllable one. By tuning a single, interpretable “risk tolerance” hyperparameter ($\lambda_{cost}$), a practitioner can explicitly guide the agent’s policy on a spectrum from aggressive and performance-focused to conservative and security-focused. The primary contribution of this work is therefore a practical, tunable, and analyzable framework for creating more intelligent and truly resilient security for decentralized AI systems.
@@ 9 Future Work
Immediate next steps could focus on expanding the capabilities of the agent. The “Defense Arsenal” could be enriched with a wider array of aggregation rules, including more sophisticated defenses against collusion (e.g., FoolsGold) or attacks specifically targeting non-IID data. Similarly, the diagnostic “State Vector” could be enhanced with more nuanced, information-rich metrics. A particularly interesting direction would be to investigate the agent’s resilience to adversaries that directly attempt to manipulate multiple state metrics simultaneously, which would test the limits of our lightweight diagnostic approach.
@@
Final accuracy and standard deviation of accuracy across the final 10 rounds. The following tables provide the final performance metrics for each aggregation strategy under the three distinct experimental scenarios. The “Final Accuracy” is the accuracy recorded at the final communication round (Round 50 or 75). “Std Dev (Last 10)” measures the standard deviation of the accuracy over the last 10 rounds, serving as a proxy for the stability of the learned model.
@@
Percentage of rounds each aggregation rule was chosen. This table details the learned policy of the FedStrategist agent under the “Stealth Attack” scenario ($\beta=0.5$) for different values of the risk-tolerance parameter, $\lambda_{cost}$. The data shows a clear shift in policy from risk-seeking (favoring the low-cost ‘FedAvg’) to risk-averse (exploring safer, more expensive rules) as $\lambda$ increases.
@@
  • Reward Function: $R_{t}=(\text{Acc}_{t}-\text{Acc}_{t-1})-(\lambda\times C_{j})$. Heuristic costs were assigned as $C_{FedAvg}=0.1$, $C_{Median}=0.4$, and $C_{Krum}=0.8$. The risk-tolerance parameter $\lambda_{cost}$ was varied in $\{0.1,0.5,1.0,2.0\}$.
@@
Complete Python source code. A zipped archive containing the complete Python source code for our custom simulation framework, ‘FedStrategist’, as well as the shell scripts used to run the experiments and the Python script to generate the plots, ensuring full reproducibility of our findings. The code is available here.
@@
All relevant data and source code are available to ensure the reproducibility of our findings. The raw CSV log files for all experimental runs presented in this study, the Python scripts used to generate the plots, and the complete source code for the ‘FedStrategist’ simulation framework are permanently archived and publicly accessible on Zenodo. The dataset can be accessed at https://doi.org/10.5281/zenodo.16068113, under DOI: 10.5281/zenodo.16068113. The development repository is also maintained on GitHub at https://github.com/rafidhaque/FedStrategist.

### v2

@@ 7.1 The Brittleness of Static Defenses and the Need for Adaptation
7.2 Interpreting the ”Paradox of the Losing Winner”
@@
Our adaptive agent, in contrast, behaved as an effective risk manager. Its diagnostic metrics correctly identified a persistent attack, and it consequently learned a conservative policy that prioritized model integrity over raw, compromised performance. The slightly lower accuracy of the adaptive agent is not a failure, but rather the measurable ”cost of security” paid to produce a reliable final model. This highlights that the objective in secure FL is not merely accuracy, but trustworthy accuracy.
@@ 7.3 The Controllable Risk Posture of FedStrategist
This demonstrates that FedStrategist is not a black box but a tunable controller. It reframes the defense problem from a search for a single ”best” algorithm to providing a system where a practitioner can define their desired security posture via a single, interpretable parameter. This tunability is a critical feature for deploying FL systems in real-world applications with varying risk requirements.
@@ 8 Conclusion
Most significantly, our analysis of the stealth attack scenario revealed a crucial insight: optimizing for raw test accuracy alone is insufficient and can be misleading. We showed that FedStrategist successfully navigates the ”paradox of the losing winner,” learning a conservative policy to guarantee model integrity even when a compromised, non-robust baseline appeared to perform better. Furthermore, we demonstrated that our framework is not a black box, but a controllable one. By tuning a single, interpretable ”risk tolerance” hyperparameter ($\lambda_{cost}$), a practitioner can explicitly guide the agent’s policy on a spectrum from aggressive and performance-focused to conservative and security-focused. The primary contribution of this work is therefore a practical, tunable, and analyzable framework for creating more intelligent and truly resilient security for decentralized AI systems.
@@ 9 Future Work
Immediate next steps could focus on expanding the capabilities of the agent. The ”Defense Arsenal” could be enriched with a wider array of aggregation rules, including more sophisticated defenses against collusion (e.g., FoolsGold) or attacks specifically targeting non-IID data. Similarly, the diagnostic ”State Vector” could be enhanced with more nuanced, information-rich metrics. A particularly interesting direction would be to investigate the agent’s resilience to adversaries that directly attempt to manipulate multiple state metrics simultaneously, which would test the limits of our lightweight diagnostic approach.
@@
Final accuracy and standard deviation of accuracy across the final 10 rounds. The following tables provide the final performance metrics for each aggregation strategy under the three distinct experimental scenarios. The ”Final Accuracy” is the accuracy recorded at the final communication round (Round 50 or 75). ”Std Dev (Last 10)” measures the standard deviation of the accuracy over the last 10 rounds, serving as a proxy for the stability of the learned model.
@@
Percentage of rounds each aggregation rule was chosen. This table details the learned policy of the FedStrategist agent under the ”Stealth Attack” scenario ($\beta=0.5$) for different values of the risk-tolerance parameter, $\lambda_{cost}$. The data shows a clear shift in policy from risk-seeking (favoring the low-cost ’FedAvg’) to risk-averse (exploring safer, more expensive rules) as $\lambda$ increases.
@@
[…]
@@
Complete Python source code. A zipped archive containing the complete Python source code for our custom simulation framework, ’FedStrategist’, as well as the shell scripts used to run the experiments and the Python script to generate the plots, ensuring full reproducibility of our findings. The code is available here.
@@
All relevant data and source code are available to ensure the reproducibility of our findings. The raw CSV log files for all experimental runs presented in this study, the Python scripts used to generate the plots, and the complete source code for the ’FedStrategist’ simulation framework are permanently archived and publicly accessible on Zenodo. The dataset can be accessed at https://doi.org/10.5281/zenodo.16068113, under DOI: 10.5281/zenodo.16068113. The development repository is also maintained on GitHub at https://github.com/rafidhaque/FedStrategist.